OpenStudy (anonymous):

please help me with this (8x^2+4x+1) + (5x^2+4x+16)

OpenStudy (anonymous):

a x b^2 + c x b^2 = (a + c) x b^2 a x b + c x b = (a + c) x b Collect like terms and add the coefficients.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @MikiaseKebede And you're welcome!

OpenStudy (anonymous):

thank you for your help and I hpe you the same @tcarroll010